XYZ Company is considering the purchase of a new piece of equipment and
has gathered the following information about the purchase:

Initial investment ..............     ?
Annual cost savings .............  $20,000
Salvage value in 6 years ........  20% of original cost of the equipment
Repair in 4 years ...............  $14,000
Cost of capital .................     10%
Life of project .................   6 years

The net present value of this new equipment was -$37,779.

Calculate the salvage value for this piece of equipment.
Answer #1

Solution:

Let Initial investment = X

Salvage value = X * 20% = 0.20 X

Now

Present value of cash inflows - Present value of cash outflows = -$37,779

($20,000 * Cumulative PV factor at 10% for 6 periods) + (0.20X * PV Factor at 10% for 6th period) - X - $14,000*PV factor at 10% for 4th period = -$37,779

($20,000 * 4.35526) + (0.20X * 0.56447) - X - ($14,000*0.68301) = -$37,779

= $87,105+ 0.11289X - X - $9,562 = -$37,779

0.8871 X = $115,322

X = $130,000

Salvage value = $130,000*20% = $26,00